OpenStudy (kayders1997):

Water is slowing at the rate of 50m^3/min from a concrete conical reservoir of the base radio us 45m and height 6m How fast is the water level falling when the water is 5 m deep? Give answer in cm/min

OpenStudy (kayders1997):

https://www.twiddla.com/2552688

OpenStudy (kayders1997):

I'll show work on what I did on there

OpenStudy (danjs):



OpenStudy (danjs):

\[V = \frac{ 1 }{ 3 }*\pi*r^2*h\]

OpenStudy (kayders1997):

I have a question

OpenStudy (kayders1997):

I called the height 6 but wouldn't that h have to be 6-h than?

OpenStudy (danjs):

Find: dh/dt Write the volume in terms of just h by using similar triangles to relate r and h. \[\frac{ 45 }{ 6 }=\frac{ r }{ h }\] \[r=\frac{ 45 }{ 6 }*h\]

OpenStudy (kayders1997):

Why wouldn't it be 6-h?

OpenStudy (kayders1997):

Oh wait

OpenStudy (danjs):

the entire height is 6m, the arbitrary height is h for the smaller triangle

OpenStudy (danjs):

Now you have a relation between r and h by similar triangles... put that r into the Area \[A = \frac{ 1 }{ 3 }*\pi*(\frac{ 45 }{ 6 }h)^2*h\] Simplify that and take the time derivative of the whole thing

OpenStudy (kayders1997):

But if It was for the large one and the small one was 6 it would be h-6 than? For that part of the height

OpenStudy (kayders1997):

I'm sorry if I'm confusing you

OpenStudy (danjs):